# Explainable post-training bias mitigation with distribution-based fairness metrics

### v2

@@ Abstract
We develop a novel optimization framework with distribution-based fairness constraints for efficiently producing demographically blind, explainable models across a wide range of fairness levels. This is accomplished through post-processing, avoiding the need for retraining. Our framework, which is based on stochastic gradient descent, can be applied to a wide range of model types, with a particular emphasis on the post-processing of gradient-boosted decision trees. Additionally, we design a broad class of interpretable global bias metrics compatible with our method by building on previous work. We empirically test our methodology on a variety of datasets and compare it to other methods.
@@ 1 Introduction
There is a comprehensive body of research on fairness metrics and bias mitigation. The bias mitigation approaches discussed in the survey paper [42] depend on the operational flow of model development processes and fall into one of three categories: pre-processing methods, in-processing methods, and post-processing methods. Pre-processing methods modify datasets before model development to reduce the bias in trained models. In-processing methods modify the model development procedure itself. Finally, post-processing methods adjust already-trained models to be less biased. Each of these categories of approach has unique benefits and drawbacks which affect their application in business settings.
@@
In this work, we build upon the ideas of [29, 65] to develop an optimization framework with distribution-based fairness constraints for producing demographically blind, explainable models across a wide range of fairness levels via post-processing. Our framework applies to various types of models, though we specifically emphasize the post-processing of gradient-boosted decision trees. Unlike neural networks, incorporating fairness constraints into these models is challenging as one must adapt the boosting process itself [49]. Our methodology supports metrics compatible with gradient descent including a wide range of metrics of interest to the financial industry (see Section 2.3); we also extend the class of global metrics discussed in [29, 44, 3].
To motivate the discussion further, consider the joint distribution $(X,Y,G)$, where $X=(X_{1},\dots,X_{n})$ is a vector of features, $Y\in\mathbb{R}$ is the response variable, and $G\in\{0,1,\dots,K-1\}$ represents the protected attribute. Given the various considerations that influence the model development in financial institutions we outline some desired properties for bias mitigation:
@@
    Explainability. Fairer models should be explainable, as regulations in FIs require applicants to be informed of factors leading to adverse credit decisions [^1] See [24] for further discussion of regulatory constraints impacting FIs, and Section 2.3 for further details on explainability.. That is, we require that an existing explainability algorithm (e.g. [52, 66, 40, 39, 41, 19, 69]) can be directly applied to, or extended for, the fairer model at a reasonably low computational cost.
@@
Many of the aforementioned bias mitigation approaches do not meet the above criteria. For example, post-processing methods that employ optimal transport [29, 36] address criteria $(i)$, $(ii)$, and $(v)$ but produce models that explicitly depend on the protected attribute (except [45], where the dependence is removed). These approaches also transform the trained model in a manner that makes explainability difficult.
@@
Model-agnostic methods, such as [50, 43], may address criteria (i)(i)-$(iv)$ but they rely on Bayesian optimization which has limited optimization power [20] , posing challenges for $(v)$. Model-specific approaches, such as [29, 65], are appealing in light of their use of distribution-based bias metrics and gradient-based techniques. However, [29] considers logistic regression models that have limited predictive capability and the computation of the gradients hinges on the optimal coupling. The method in [65] considers neural networks for ROC-based fairness constraints, which are natural candidates for gradient-based methods, but these types of models are known to underperform on tabular data compared to tree ensembles [59]. In these cases, model flexibility $(ii)$ and explainability $(iv)$ may be compromised. Finally, [51] proposes an extension of XGBoost that integrates fairness constraints into its gradient and Hessian computations. These constraints are required to be in the form of expected losses, making them incompatible with most integral probability metrics such as Wasserstein that lack per-sample loss formulations.
Overall, both pre-processing and in-processing approaches often require costly model re-training in order to achieve fairer models across varying levels of bias (i.e., the efficient bias-performance frontier), which can make model development prohibitively expensive, especially when datasets are large.
@@
In this work, we address all criteria (i)(i)$(v)$ by proposing a novel post-processing approach to bias mitigation. Specifically, we introduce a broad family of global bias metrics and develop several estimators suitable for optimization via gradient descent. Next, we construct several parameterized families of post-processed models that can be adapted from any trained, demographically blind, and explainable model architecture. These models are designed to remain explainable and differentiable with respect to their parameters. Finally, we identify bias-performance efficient frontiers of these constructed families by optimizing under fairness constraints, expressed through our proposed global bias estimators. This optimization is carried out using stochastic gradient descent, in the same manner that one may train a neural network. We elaborate on the most salient technical details of this strategy below.
To address criterion $(i)$, when $G\in\{0,1\}$ is binary, we propose a broad class of distribution-based bias metrics of the form
@@
where $c(\cdot,\cdot)$ is a cost function, $F_{f(X)|G=k}$ is the cumulative distribution function of $f(X)|G=k$, and $\mu$ is a probability measure signifying the importance of the classifier associated with threshold $t\in\mathbb{R}$ which may in principle, depend on $(f,P_{G},P_{X|G})$ as in [65, 3]. For a raw probability score, $f$ in (1.1) is replaced with $\operatorname{logit}(f)$. This formulation encompasses many metrics including the 1-Wasserstein metric, the energy distance [62], and others [3, 65]. It may also be generalized to non-binary protected attributes as in [29, 44]. To design estimators for these global metrics, we employ relaxation methods in the spirit of those in [65] and, to justify their use in numerical optimization, perform an asymptotic analysis; see Section 3.
We may now design generic model-flexible methods for producing families of post-processed models. Given a trained regressor or raw probability score model $f_{*}$, we pick a vector $w=(1,w_{1},\dots,w_{m})(x;f_{*})$ of weight functions (or encoders) and construct the family of models:
@@
where $\theta\in\mathbb{R}^{m+1}$ is learnable, and $w$ may generally depend on the model representation; see Section 4.2. Typically, in view of criteria $(iii)$ and $(iv)$, we require that $f_{*}$ be demographically blind and explainable, and that $w$ share these properties.
@@
Crucially, the above minimization problem is linear in $w$. Thus, in contrast to the post-processed model families in [43], this setup circumvents the lack of differentiability of the trained model. This enables the use of gradient-based methods as opposed to Bayesian search even when $f_{*}$ is discontinuous (e.g., tree-based ensembles). Consequently, we may efficiently post-process any model while optimizing a high-dimensional parameter space with stochastic gradient descent.
@@
Clearly, the selection of the encoders is crucial for ensuring the demographic blindness and explainability of post-processed models. While they can be constructed in various ways, we present three particular approaches that yield families of explainable models where the encoders are selected in the form of additive models, weak learners (for tree ensembles), and finally, explanations; see Section 4.
Our approach quickly generates demographically blind, explainable models with strong bias-performance trade-offs. We empirically compare it to [43] as well as an explainable optimal transport projection method based on [45] across various datasets [4, 47, 2]. We also discuss how dataset properties impact performance and propose strategies to address overfitting.
@@ 2.1 Notation and hypotheses
Given a model $f$ and features $X$, we set $Z:=f(X)$ and denote model subpopulations as $Z_{k}:=f(X)|G=k$, $k\in\mathcal{G}$. The subpopulation cumulative distribution function (CDF) of $Z_{k}$ is denoted by $F_{k}(t):=F_{f(X)|G=k}(t)=\mathbb{P}(f(X)\leq t|G=k)$, and the corresponding generalized inverse (or quantile function) $F_{k}^{[-1]}$ is defined by $F_{k}^{[-1]}(p):=F_{f(X)|G=k}^{[-1]}(p)=\inf_{x\in\mathbb{R}}\big{\{}p\leq F_{k}(x)\big{\}}$, for each $k\in{\cal G}$. Finally, a derived classifier $f_{t}(x;f)$ associated with the model $f$ and a threshold $t\in\mathbb{R}$ is defined by $f_{t}(x;f)=\mathbbm{1}_{\{f(x)>t\}}$.
@@ 2.2 Classifier fairness definitions and biases
Numerous works have investigated statistical parity [31, 18, 22, 29] and equal opportunity [32, 65] fairness criteria. Meanwhile, the $\mathcal{A}$-based parity may be viewed as a generalization of statistical parity, equalized odds, and equal opportunity biases. For example, letting $\mathcal{A}=\{\Omega\}$ produces the statistical parity criterion, letting $\mathcal{A}=\{\{Y=0\},\{Y=1\}\}$ produces the equalized odds criterion, and letting $\mathcal{A}=\{\{Y=1\}\}$ produces the equal opportunity criterion. It may also be viewed as an extension of conditional statistical parity in [63] where the true response variable $Y$ is treated as a factor in determining fairness. The methods introduced in this work may be adapted to any $\mathcal{A}$-based parity criterion, but we focus on statistical parity (i.e., where $\mathcal{A}=\{\Omega\}$) for simplicity. We now present the definition of the classifier bias for statistical parity.
@@ 2.3 Distribution-based fairness metrics
Definition 2.4 (Wasserstein model bias [44]).
@@
According to [3], when scores have continuous distributions, (2.4) is equal to $W_{1}(P_{F_{Z}(Z_{0})},P_{F_{Z}(Z_{1})})$, where $Z_{k}$, $k\in\{0,1\}$ are subpopulation scores. When $P_{Z}$ has atoms, the above relationship generally does not hold (see Example E.1). Nevertheless, it can be generalized. Specifically, we have the following result.
@@
Proposition 2.2 remains true if $c(x,y)=d(x,y)^{p}$, where $d(\cdot,\cdot)$ is a metric, with $p\geq 1$. In that case, $\mathscr{T}_{c}({F_{0}}_{\#}\mu,{F_{1}}_{\#}\mu)^{1/p}=W_{p}({F_{0}}_{\#}\mu,{F_{1}}_{\#}\mu;d)$.
@@ 3 Bias metrics approximations for stochastic gradient descent
However, due to the presence of indicator functions, $\hat{F}_{k}(t,\theta)$ is in general not differentiable in $\theta$. Thus, substituting $F_{k}$ for $\hat{F}_{k}$ in (3.1) may result in bias metrics that are not differentiable. To address this issue, we adopt a relaxation of the formulation (3.1), inspired by [65], which enables the construction of various differentiable approximations suited to stochastic gradient descent.
@@ 3.1 Relaxation approximation
By Lemma D.1, $F^{(s)}_{k}$ is a globally Lipschitz CDF approximating $F_{k}$, where $\lim_{s\to\infty}F_{k}^{(s)}(t)=F_{k}(t)$, $\forall t\in\mathbb{R}$, and
@@ 3.2 Bias estimators
Following the discussion above, we propose several approaches to estimate the relaxed bias metric
@@
employing the estimator (3.3). Throughout, we assume that ${\rm Lip}(r_{s})\leq s$ for every $s\in\mathbb{R}_{+}$. We also assume that $\mu$ is fixed. We extend the estimators introduced below to cases where $\mu$ depends on $(f,P_{G},P_{X|G})$, such as the metrics in [65, 3], in Appendix F.
@@ 3.2.2 Threshold-discrete estimator
Finally, the estimator (3.7) and the proof of Theorem 3.2 can be easily adapted to cases where the density $\rho$ has finite support or decays exponentially fast.
@@
where $\hat{P}_{f(X;\theta)|G=k}$, $k\in\{0,1\}$, denotes the empirical distribution of the model’s scores. The gradient of the right-hand side in (3.9) can be computed explicitly using the optimal transport coupling; see [29, Lemma 3].
@@ 4.2 Explainable bias mitigation through output perturbation
where $\theta\in\Theta$ is a learnable parameter [^4] For calibration purposes, one can introduce an additional parameter $\alpha$ in (4.1), yielding models of the form $\alpha f_{*}-\theta\cdot w(x;f_{*})$, where $f_{*}$ is the base model..
@@
For instance, when $E=\varphi^{\text{\tiny\it ME}}$, relation (4.3) holds. In this case, it is essential to use explanation algorithms that efficiently compute the marginal Shapley value for both the trained model and the encoders. For example, if $f_{*}$ is a tree ensemble and $T$, $L$ and $D$ denote its number of trees, maximum number of leaves, and background dataset respectively, the per-instance complexity of explanations is $O(T\cdot L\cdot|D|)$ for interventional TreeSHAP [41] and $O(T\cdot\log{L})$ for the algorithm of [19] for oblivious trees. The dependence on $|D|$ in the former but not the latter can make a big difference on practical feasibility for large-scale applications. For other cases, model-agnostic approximations (e.g., KernelSHAP [39] or Monte Carlo methods [66]) can be employed.

### v3

@@ Abstract
We develop a novel bias mitigation framework with distribution-based fairness constraints suitable for producing demographically blind and explainable machine-learning models across a wide range of fairness levels. This is accomplished through post-processing, allowing fairer models to be generated efficiently without retraining the underlying model. Our framework, which is based on stochastic gradient descent, can be applied to a wide range of model types, with a particular emphasis on the post-processing of gradient-boosted decision trees. Additionally, we design a broad family of global fairness metrics, along with differentiable and consistent estimators compatible with our framework, building on previous work. We empirically test our methodology on a variety of datasets and compare it with alternative post-processing approaches, including Bayesian search, optimal transport projection, and direct neural network training.
@@ 1 Introduction
In this work, we propose a framework that addresses the problem of ML model bias in the financial industry setting under regulatory constraints. In such contexts, bias mitigation methods must satisfy multiple criteria: they must accommodate diverse predictive models, avoid direct use of the protected attribute, yield models that maintain explainability, and be computationally efficient to enable the construction of models along a fairness–performance frontier. Furthermore, it is often crucial to employ global bias metrics that evaluate fairness over the entire score distribution, since the decision thresholds of classifiers are not typically known at the training stage and may differ across downstream applications. To accomplish this, we employ global bias metrics such as those used in [3, 4, 5] and gradient descent based optimization techniques as in [6, 7].
There is a comprehensive body of research on fairness metrics and bias mitigation. The bias mitigation approaches discussed in the survey paper [8] depend on the operational flow of model development processes and fall into one of three categories: pre-processing methods, in-processing methods, and post-processing methods. Pre-processing methods modify datasets before model development to reduce the bias in trained models. In-processing methods modify the model development procedure itself. Finally, post-processing methods adjust already-trained models to be less biased. Each category presents trade-offs that affect their suitability for real-world applications, particularly in regulated domains such as finance.
@@
While these approaches are often useful, their application in financial services may be limited by the aforementioned practical and regulatory considerations. To elaborate, we seek a bias mitigation strategy that meets the following criteria:
@@
    Explainability. Fairer models should be explainable, as regulations in FIs require applicants to be informed of factors leading to adverse credit decisions [^1] See [26] for further discussion of regulatory constraints impacting FIs, and Section 2.3 for further details on explainability.. That is, we require that an existing explainability algorithm (e.g., [27, 28, 29, 30, 31, 32, 33]) can be applied to or extended for the fairer model at a reasonable computational cost.
@@
For many existing in-processing bias mitigation approaches, these criteria pose challenges. Some notable methods, such as [7, 6], are appealing in light of their use of distribution-based bias metrics and gradient-based optimization. Specifically, [7] trains logistic regression models with a loss penalized by the 1-Wasserstein barycenter of subpopulation score distributions, while [6] introduce a custom loss integrating a ROC-based fairness metric into neural network training. However, they limit model flexibility $(ii)$ and explainability $(iv)$, and they preclude the use of tree ensembles, which are often preferred for tabular data [34]. Similarly, [35] proposes an extension of XGBoost that integrates fairness constraints into its gradient and Hessian computations. These constraints must be expressed as expected losses, making them incompatible with most integral probability metrics such as Wasserstein that lack per-sample loss formulations.
Some in-processing methods avoid reliance on specific model architectures, improving flexibility. For example, [14] introduces a model-agnostic approach that incorporates fairness into hyperparameter tuning via Bayesian optimization. While this can satisfy criteria (i)(i)$(iv)$, the reliance on Bayesian search limits optimization power [36], posing challenges for criterion $(v)$.
Post-processing methods such as [21] also rely on Bayesian optimization and therefore have similar limitations in computational efficiency. Other post-processing methods, such as [7] and [3], employ optimal transport to align score distributions across subpopulations. These methods address criteria $(i)$, $(ii)$, and $(v)$, but produce models that explicitly depend on the protected attribute – an exception being [37], where the dependence is removed. Furthermore, these approaches transform the trained model in ways that hinder explainability.
Overall, existing pre-processing and in-processing approaches often require costly retraining to obtain fairer models across different bias levels (i.e., the efficient bias-performance frontier), which may be computationally impractical for large datasets. Additionally, such approaches reduce flexibility, as the fairness metric is embedded into the training process. This makes later adaptation to regulatory or institutional changes difficult. Conducting bias mitigation in the post-processing stage, by contrast, allows practitioners to select or refine the fairness metric after model training, providing flexibility and adaptability.
@@
Our novel post-processing framework for bias mitigation, motivated by [7, 6], addresses all criteria (i)(i)$(v)$. We introduce a broad family of global bias metrics and develop differentiable estimators suitable for optimization via gradient descent. We then construct parameterized families of post-processed models that can be adapted from any trained, demographically blind, and explainable model architecture. These model families are designed to remain explainable and allow efficient gradient-based optimization. Finally, we construct bias-performance efficient frontiers of these model families by optimizing under fairness constraints, expressed through our proposed global bias estimators. This optimization is carried out using stochastic gradient descent, in the same manner that one may train a neural network.
Unlike in-processing approaches, which must embed the fairness metric into the training process, our method performs fairness optimization over a family of correctors (encoders), enabling fast and flexible construction of fair models. This is particularly well suited for tree-based models, such as gradient-boosted decision trees, where fairness constraints are hard to incorporate during training. We elaborate on the most salient technical details of this strategy below.
To ground the discussion, we consider the joint distribution $(X,Y,G)$, where $X=(X_{1},\dots,X_{n})$ is a vector of features, $Y\in\mathbb{R}$ is the response variable, and $G\in\{0,1,\dots,K-1\}$ represents the protected attribute. Addressing criterion $(i)$, we propose a broad class of distribution-based bias metrics that quantify disparities between the distributions of model scores across protected groups. When $G\in\{0,1\}$ is binary, the metric takes the form
@@
where $c(\cdot,\cdot)$ is a cost function, $F_{f(X)|G=k}$ is the cumulative distribution function of $f(X)|G=k$, and $\mu$ is a probability measure signifying the importance of the classifier associated with threshold $t\in\mathbb{R}$ which may in principle, depend on $(f,P_{G},P_{X|G})$ as in [6, 24]. For a raw probability score, $f$ in (1.1) is replaced with $\operatorname{logit}(f)$. This formulation encompasses many metrics including the 1-Wasserstein metric, the energy distance [38], and others [24, 6]. Moreover, it generalizes naturally to non-binary protected attributes as in [7, 20]. We design estimators for these metrics via relaxation methods inspired by [6] and perform an asymptotic analysis to justify their use in numerical optimization; see Section 3.
We can now design generic, model-flexible methods to produce families of post-processed models. Given a trained regressor or raw probability score model $f_{*}$, we select a vector $w=(1,w_{1},\dots,w_{m})(x;f_{*})$ of weight functions (or encoders) and construct the corresponding family of models:
@@
where $\theta\in\mathbb{R}^{m+1}$ is learnable, and $w$ may generally depend on the underlying model representation; see Section 4.2. In accordance with criteria $(iii)$ and $(iv)$, we typically require that both $f_{*}$ and $w$ be demographically blind and explainable.
Finally, we tackle criterion $(v)$ and seek models in ${\cal F}(f_{*},w)$ whose bias-performance trade-off is optimal – that is, the least biased among similarly performing models. To construct the efficient frontier of ${\cal F}(f_{*};w)$, adapting the approaches in [7, 6], we solve a minimization problem with a fairness penalization [39, 40]:
$\theta^{*}(\omega):={\rm argmin}_{\theta}\{{\cal L}(f_{\theta})+\omega\mathcal{B}(f_{\theta}|X,G)\},$
@@
Crucially, the above minimization problem is linear in $w$. Unlike the post-processing approach in [21], this setup circumvents the lack of differentiability of the trained model, enabling the use of gradient-based methods as opposed to Bayesian search even when $f_{*}$ is discontinuous (e.g., tree-based ensembles). Consequently, we may efficiently post-process any model while optimizing a high-dimensional parameter space via stochastic gradient descent.
@@
Clearly, the choice of encoder family is key to ensuring both demographic blindness and explainability in post-processed models. We propose three such families – based on additive models, weak learners (for tree ensembles), and model explanations – that yield interpretable and flexible corrections; see Section 4.
Our approach enables fast construction of demographically blind, explainable models with favorable bias-performance trade-offs. We evaluate it on both synthetic and real-world datasets [41, 42, 43], comparing against the post-processing method of [21] and an in-processing neural network approach adapted from [6]. We find that our post-processing method yields strong bias-performance frontiers. We also examine how dataset characteristics influence outcomes and suggest strategies to mitigate overfitting.
@@ 2.1 Notation and hypotheses
Given a model $f$ and features $X$, we set $Z:=f(X)$ and denote model subpopulations as $Z_{k}:=f(X)|G=k$, $k\in\mathcal{G}$. The subpopulation cumulative distribution function (CDF) of $Z_{k}$ is denoted by
$F_{k}(t):=F_{f(X)|G=k}(t)=\mathbb{P}(f(X)\leq t|G=k),$
and the corresponding generalized inverse (or quantile function) $F_{k}^{[-1]}$ is defined by
$F_{k}^{[-1]}(p):=F_{f(X)|G=k}^{[-1]}(p)=\inf_{x\in\mathbb{R}}\big\{p\leq F_{k}(x)\big\},$
for each $k\in{\cal G}$. Finally, given a threshold $t\in\mathbb{R}$, the corresponding binary classifier is defined as $f_{t}(x;f)=\mathbbm{1}_{\{f(x)>t\}}$.
@@ 2.2 Classifier fairness definitions and biases
Numerous works have investigated statistical parity [13, 10, 9, 7] and equal opportunity [15, 6] fairness criteria. Meanwhile, the $\mathcal{A}$-based parity may be viewed as a generalization of statistical parity, equalized odds, and equal opportunity biases. For example, letting $\mathcal{A}=\{\Omega\}$ produces the statistical parity criterion, letting $\mathcal{A}=\{\{Y=0\},\{Y=1\}\}$ produces the equalized odds criterion, and letting $\mathcal{A}=\{\{Y=1\}\}$ produces the equal opportunity criterion. It may also be viewed as an extension of conditional statistical parity in [44] where the true response variable $Y$ is treated as a factor in determining fairness.
The methods introduced in this work may be adapted to any $\mathcal{A}$-based parity criterion, but we focus on statistical parity (i.e., where $\mathcal{A}=\{\Omega\}$) for simplicity. We now present the definition of the classifier bias for statistical parity.
@@ 2.3 Distribution-based fairness metrics
Definition 2.4 (Wasserstein-1 model bias [20]).
@@
According to [24], when the score distributions are continuous, (2.4) equals $W_{1}(P_{F_{Z}(Z_{0})},P_{F_{Z}(Z_{1})})$, where $Z_{k}$, $k\in\{0,1\}$, are subpopulation scores. When $P_{Z}$ has atoms, the above relationship generally does not hold (see Example E.1). Nevertheless, it can be generalized. Specifically, we have the following result.
@@
Proposition 2.2 remains valid if $c(x,y)=d(x,y)^{p}$, where $d(\cdot,\cdot)$ is a metric, with $p\geq 1$. In that case, we have
$\mathscr{T}_{c}({F_{0}}_{\#}\mu,{F_{1}}_{\#}\mu)^{1/p}=W_{p}({F_{0}}_{\#}\mu,{F_{1}}_{\#}\mu;d).$
@@ 3 Bias metrics approximations for stochastic gradient descent
However, due to the presence of indicator functions, $\hat{F}_{k}(t{\color[rgb]{.75,.5,.25};}\theta)$ is in general not differentiable in $\theta$. Thus, substituting $F_{k}$ for $\hat{F}_{k}$ in (3.1) may result in bias metrics that are not differentiable. To address this issue, we adopt a relaxation of the formulation (3.1), inspired by [6], which enables the construction of various differentiable approximations suited to stochastic gradient descent.
@@ 3.1 Relaxation approximation
By Lemma D.1, $F^{(s)}_{k}$ is a globally Lipschitz CDF approximating $F_{k}$, where $\lim_{s\to\infty}F_{k}^{(s)}(t;\theta)=F_{k}(t;\theta)$, $\forall t\in\mathbb{R}$, and
@@ 3.2 Bias estimators
Following the discussion above, we propose several methods to estimate the relaxed bias metric
@@
employing the estimator (3.3). Throughout, we assume that ${\rm Lip}(r_{s})\leq s$ for all $s\in\mathbb{R}_{+}$, and that $\mu$ is fixed. We extend the estimators introduced below to cases where $\mu$ depends on $(f,P_{G},P_{X|G})$, such as the metrics in [6, 24], in Appendix F.
@@ 3.2.2 Threshold-discrete estimator
Finally, note that the estimator (3.7) and the proof of Theorem 3.2 can be readily adapted to cases where the density $\rho$ has finite support or decays exponentially fast.
@@
where $\hat{P}_{f(X;\theta)|G=k}$, $k\in\{0,1\}$, denotes the empirical distribution of the model’s scores. The gradient of the right-hand side in (3.9) can then be computed explicitly using the optimal transport coupling; see [7, Lemma 3].
@@ 4.2 Explainable bias mitigation through output perturbation
where $\theta\in\Theta$ is a learnable parameter [^4] For calibration purposes, one can introduce an additional parameter $\alpha$ in (4.1), yielding models of the form $\alpha f_{*}-\theta\cdot w(x;f_{*})$..
@@
For instance, when $E=\varphi^{\text{\tiny\it ME}}$, relation (4.3) holds. In this case, it is essential to use explanation algorithms that efficiently compute the marginal Shapley value for both the trained model and the encoders.
For example, if $f_{*}$ is a tree ensemble and $T$, $L$ and $D$ denote its number of trees, maximum number of leaves, and background dataset respectively, the per-instance complexity of explanations is $O(T\cdot L\cdot|D|)$ for interventional TreeSHAP [31] and $O(T\cdot\log{L})$ for the algorithm of [32] for oblivious trees. The dependence on $|D|$ in the former but not the latter can make a big difference on practical feasibility for large-scale applications. For other cases, model-agnostic approximations (e.g., KernelSHAP [30] or Monte Carlo methods [28]) can be employed.